OpenStudy (anonymous):

f(x) = 9x-4ln(x) X>0 I get the critical point of 4/9... For some reason I'm screwing up finding the interval where x is increasing and decreasing. I'm tired! Please help.

OpenStudy (anonymous):

When testing x values do I use f(x) or f'(x)?

OpenStudy (anonymous):

f'(x) > 0 means f(x) is increasing

OpenStudy (anonymous):

And then test to the right and left of 4/9, right?

OpenStudy (anonymous):

i draw a number line and mark the zeros. then i check the intervals... sign check is sufficient since you only need + or -

OpenStudy (anonymous):

Can you check if my deriv is wrong? I get f'(x )= 9-4/x...

OpenStudy (anonymous):

I know the answer is that it is decreasing from (0,4/9) but I can't get it to be a negative number in my calculations lol

OpenStudy (anonymous):

f'(x) = 9 - 4/x = 0 => x = 4/9

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay! Whoops I see my error! Thanks a lot, I was testing with numbers lower than 0, whoops!

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

you're welcome